# An adaptive finite element multigrid solver using GPU acceleration

### v2

@@ 3 Data-parallel accelerators
Different libraries have been proposed to aid with the development of GPU-accelerated software. OpenACC and OpenMP provide pragma-based access to the compute power of GPUs that is very simple to use but also limited in the potential for optimizations. SYCL [39] and Kokkos [41] provide a middle-ground with more complex usage but also more flexibility. Machine learning libraries such as PyTorch [31] and jax [11] follow a new paradigm for GPU programming where the application code is specified in a high-level language and a backend, such as XLA [36] or ATen [31], generates highly efficient, hardware-specific code. This concept is sometimes also referred to as domain-specific language, e.g. [25]. Recently, Triton [40] was proposed as a complement to existing machine learning libraries. It still allows for highly simplified programming of accelerators compared to CUDA and ROCm but provides more control than PyTorch and jax, and hence typically yields more efficient code.
For specific applications, such as linear algebra or discrete Fourier transform, also highly optimized libraries exist. These provide very high performance while, through the targeted use case, also allow for much simpler usage than, e.g., CUDA. In our work, we will build on a linear algebra formulation of our adaptive PDE solvers. This allows us to use the cuBLAS [28] and cuSPARSE libraries [29] that provide high performance while being simple to use with a software interface that is roughly comparable to those of the conventional BLAS library. Furthermore, cuBLAS and cuSPARSE can still be combined with native CUDA code when necessary, i.e. when an operation cannot be expressed efficiently in linear algebra.
@@ 4 CUDA multigrid
In this section we discuss our extension of the adaptive finite element library Gascoigne 3d to use GPU accelerators for a wide range of computations. The extensions have been designed with two use cases in mind: first, non-stationary but linear systems of partial differential equations, where a single system matrix is reused in every time step or at least in many time steps, and, second, nonlinear problems that can be formulated in terms of fixed matrices that do not need re-assembly. The second case includes a pressure-projection based solvers for the nonstationary nonlinear Navier-Stokes equations but can be extended to more complex flow problems such as the Boussinesq approximation [12]. Both problem classes have in common that matrices are assembled only once and the complete workflow can be formulated in terms of matrix vector products and the solution of linear systems of equations. Algorithms 2 and 3 show the typical workflow for the problem types. Computations highlighted in blue are completely performed on the GPU, orange marks transfer of data between GPU and CPU.
@@ 4       Assemble Right Hand Side;
It is no fundamental issue to also port the computations of general nonlinear problems to a GPU. However, keeping the flexibility and support of different discretizations and triangulations is not as easily handled for the ill-structured setup of typical finite element assemblies of system matrices and residuals in the general nonlinear case. We hence defer this to future work.
@@ 4.1 General architecture
Our principal approach for the CUDA parallelization of Gascoigne 3d is to formulate all operations in terms of (potentially sparse) matrix-vector and matrix-matrix products and to realize these with the highly optimized cuBLAS [28] and cuSPARSE [29] libraries. This approach differs from the existing implementation in Gascoigne 3d where many operations such as the handling of hanging nodes or the multigrid mesh transfer are based on the local connectivity of the degrees of freedom as function of mesh and discretization. However, this functionality can easily be realized as matrix-vector products, including for the handling of hanging nodes and all mesh transfer operations. This is, e.g., the approach taken from the outset in the deal.II finite element library [2].
All high level operations in Gascoigne 3d are performed on abstract interface classes to vectors and matrices to allow for MPI parallelization [23, 5]. This has been retained in the CUDA-version so that the high-level control flow and complex algorithms (such as Newton, multigrid, GMRES) are unchanged since they only operate on the interfaces. The approach minimizes the changes that were overall required and allows to flexibly retain both the CPU and the GPU backend.
@@ 4.1.1 Storage
Gascoigne 3d uses a block-wise memory concept in which the various components of a partial differential equation (e.g. the 3 deformation unknowns in 3d solid mechanics or the pressure and the three velocity components in fluid mechanics) are stored next to each other. Matrices are double-indexed and the outer index $(i,j)$ refers to the mesh node yielding the entry $a_{i,j}\in\mathbb{R}^{n_{c}\times n_{c}}$ that stores (aligned in memory) the local couplings between the solution components as a matrix. In general, this setup helps to efficiently use caching [8, 23]. Using cuSPARSE this block-matrix format is not supported. While vectors are directly transferred to the GPU using the same memory layout, the inner matrix blocks must be resolved and matrices are stored in the usual CSR format. Fig. 2 shows a visualization of the matrix storage concept on the CPU (left) and the GPU (right).
@@ 4.1.4 Geometric Multigrid CUDA
As a concrete example, we describe in the following how geometric multi-grid algorithm in Algo. 1 is realized in CUDA-Gascoigne-3d. All computations are entirely performed on the GPU. Hence, data transfer overhead is incurred only at the beginning and end of the computations, see also Algorithms 2 and 3.
@@ 4.1.5 GMRES as linear solver on the GPU
For reasons of numerical stability, e.g. on non-uniform meshes, the multigrid iteration can often not be used directly as linear solver. They are then employed as preconditioner in a GMRES iteration. Our GMRES solver follows the approach described in [35, Section 6.5.3], using the modified Gram Schmidt algorithm for orthogonalization and Givens rotation for solving the resulting overdetermined linear system. Due to the multigrid preconditioning, we avoid stability issues in the Gram Schmidt orthogonalization and never have to perform more than 5-10 GMRES steps. The Gram Schmidt iteration is also completely run on the GPU using cuSPARSE and cuBLAS operations. As we use a highly efficient geometric multigrid solver as preconditioner, a maximum of $n_{G}\leq 10$ GMRES steps is usually required. Hence, the resulting overdetermined system for finding the GMRES solution is very small and its step-wise transformation to a diagonal matrix with Givens rotations is performed on the CPU. The cost for transferring these $n_{G}^{2}$ entries between GPU and CPU is small.. Details on GMRES are given in [35].
@@ 5 Results
All CPU computations are carried out on an AMD EPYC 7773X. For the relatively small problem sizes we consider, our scaling test in Fig. 3 shows that there is only a minor benefit from using a large number of threads. The calculations are dominated by sparse matrix-vector multiplications and the efficiency is therefore limited by the memory bandwidth. This has already been shown in a detailed parallel analysis of fluid-structure interactions based on Gascoigne [14]. A transition to matrix-free finite elements—see, e.g., [24]—would thus offer the greatest potential for improving parallel efficiency. However, this would require a fundamental redesign of the current implementation. We therefore limit the OpenMP based parallel code to 8 threads for the comparisons. The GPU computations are performed on the same machine using an NVIDIA H100 PCIe card with 80 GB of VRAM. When using the GPU to accelerate the multigrid solver, those parts of Gascoigne that reside on the CPU still run on 8 parallel threads. We always indicate wall-clock times and separately specify the different contributions of Algorithms 2 and 3.
@@ 5.1 Transport-diffusion equation
Figure 4: Transport-diffusion equation: Wall-clock times CPU vs. GPU on the finest mesh level with 1 000 00010000001\,000\,0001 000 000 degrees of freedom (left: linear scale, right: log scale).
@@
We consider the backward Euler method with a time-step size $\Delta t=0.02$. Uniform quadrilateral spatial meshes with a range of sizes given by $h=2^{-n}$ where $n\in\{4,\dots,10\}$ are used to demonstrate the scaling efficiency of the GPU accelerated version. Hence, on the coarsest mesh the problem has $6414^{7}+2^{7+1}+1=16,$ 641 degrees of freedom whereas the finest mesh comprises $0004^{10}+2^{10+1}+1\approx 1,000,$ 000 degrees of freedom. In each of the $T/\Delta t=100$ time steps, a linear system must be solved. To ensure the robustness of the method, we use the GMRES solver, preconditioned with multigrid, to approximate the linear problems, cf. Sec. 4.1.5. Fig. 4 shows the wall-clock times comparing the CPU and GPU implementations on the finest mesh level. The labels init, rhs and solve refer to the initialization (mainly assembly of system matrix), to the computation of the right hand side, and to the actual solution, respectively. Matrix assembly and the computation of the right hand side involve numerical quadrature over the mesh elements which is not ported to the GPU. Hence, no speedup is observed, GPU timings can even be higher, as copying the matrix to the GPU involves some overhead, see also Fig. 5. The time for solving the linear problems, however, is reduced from $304.75\,\mathrm{s}$ on the CPU to $9.0\,\mathrm{s}$ on the GPU, a factor of about 34; see also Table 1 for the raw data. The geometric multigrid solver is robust on all meshes such that the number of linear steps varies only slightly between the mesh levels. For large problems, the non-accelerated assembly of the right hand side rhs hence becomes dominant. In Section 5.3 we describe how this term can be efficiently transferred to the GPU by using mass lumping.
@@ 5.2 Linear elasticity
For time discretization of (6)-(7) the backward Euler method is used with a step size of $\Delta t=0.025$ resulting in $1002.5/\Delta t=$ 100 time steps.
@@ 5.3 Navier-Stokes equations
$(\mbox{div}\,\boldsymbol{u},\xi)\;=\;$ 00 0$\qquad$ $\forall\xi\in Q.$ (11)
@@
$\frac{1}{k}(\boldsymbol{u}_{h}^{m},\boldsymbol{\chi})_{*}$ $\;=\;(\boldsymbol{f}^{m-1}+\frac{1}{k}\boldsymbol{u}^{m-1}_{h},\boldsymbol{\chi})_{*}-A((\boldsymbol{u}^{m-1}_{h},p^{m-1}_{h}+q^{m-1}_{h});(\boldsymbol{\chi},0))\quad\forall\boldsymbol{\chi}\in\boldsymbol{V}_{h}.$
@@
$(\nabla q^{m}_{h},\nabla\varphi)$ $\;=\;-\frac{1}{k}(\mbox{div}\,\boldsymbol{u}_{h}^{m},\varphi)\quad\forall\varphi\in S_{h}.$
@@
$(p_{h}^{m},\varphi)_{*}$ $\;=\;(p_{h}^{m-1}+q_{h}^{m},\varphi)_{*}-(\nu\,\mbox{div}\,\,\boldsymbol{u}_{h}^{m},\varphi)\quad\forall\varphi\in Q_{h}.$
@@
Moreover, we apply a mass lumping approach also in the nonlinear term, which enables us to implement the convective term as a matrix-vector product with a pre-assembled sparse matrix, see also [21]. In particular, let $\mathcal{I}_{h}:\boldsymbol{V}\to\boldsymbol{V}_{h}$ be the nodal interpolation operator. For each $\boldsymbol{v}_{h},\;\boldsymbol{\chi}_{h}\in\boldsymbol{V}_{h}$, we approximate
@@
The error caused by mass lumping in the zeroth-order term (12) and the convective term (13) are of the same order as the polynomial approximation error and therefore do not affect the error asymptotics. This approach is sometimes denoted as ‘fully practical finite element method’ [3] as it allows for error estimates that reflect the full error including numerical quadrature. We have recently published the complete error analysis for an explicit pressure correction method based on this approximation [22]. Both theoretical analysis and numerical examples show that the explicit treatment based on the interpolation of the nonlinearity does not lead to any additional error.
@@
$-\big{(}\boldsymbol{u}_{h}\otimes\boldsymbol{u}_{j},\nabla\boldsymbol{\chi}\big{)}_{i=1}^{N_{nodes}}\approx-C_{h,x}\boldsymbol{v}^{1}-C_{h,x}\boldsymbol{v}^{2}-C_{h,x}\boldsymbol{v}^{3}.$ (15)
@@ 5.3.1 Explicit projection solver
For justification and well-posedness of this equation, see [20]. Moreover, note that the velocity solutions $\boldsymbol{u}_{h}^{m}$ are so-called predictor solutions. The corrector velocities were eliminated as described in [18]. The complete error analysis for this fully explicit variant of a pressure correction scheme is found in [22] and its extension to the Boussinesq approximation is described in [13].
@@ 5.3.3 Numerical Results
[…]
@@ 5.4 Limitations
For speed, the bottleneck of the GPU implementation is the memory bandwidth. The increase in bandwidth by a factor of 10 from the DDR4 RAM accessed by the CPU to the HBMe2 memory utilized on the H100 GPU (2 TB/s) is complemented by a similar increase in potential compute throughput. In the Navier-Stokes example, roughly $2/3$ of the GPU cycles are spend on sparse matrix-vector products, which have an arithmetic intensity of 0.39FLOP/byte (0.15FLOP/byte for the transposed variant) and therefore achieve just $\frac{1}{32}$ ($\frac{1}{83}$) of the peak double-precision performance. As part of cuSPARSE, the kernels are already highly optimized with a memory throughput of $91\%$ ($82\%$) and the low arithmetic intensity is inherent to the computational problem. Major improvements would therefore require fusing some of the roughly 3500 kernel calls needed for a single time-step together in order to reuse the data. This is not practical with the current approach that heavily leverages the cuSPARSE library. All the custom kernels together use less than $3\%$ of the GPU time. While some of these could be eliminated by assembling rectangular block matrices in Gascoigne 3D, the speedup would therefore be negligible. The most promising direction for further optimizations is therefore a switch to single precision, since this effectively doubles the arithmetic intensity.

### v3

@@ 3 Data-parallel accelerators
Different libraries have been proposed to aid with the development of GPU-accelerated software. OpenACC and OpenMP provide pragma-based access to the compute power of GPUs that is very simple to use but also limited in the potential for optimizations. SYCL [syclmain] and Kokkos [Trott2022] provide a middle-ground with more complex usage but also more flexibility. Machine learning libraries such as PyTorch [paszke2017automatic] and jax [jax2018github] follow a new paradigm for GPU programming where the application code is specified in a high-level language and a backend, such as XLA [Sabne2020_XLA] or ATen [paszke2017automatic], generates highly efficient, hardware-specific code. This concept is sometimes also referred to as domain-specific language, e.g. [Lawrence2018]. Recently, Triton [Tillet2019] was proposed as a complement to existing machine learning libraries. It still allows for highly simplified programming of accelerators compared to CUDA and ROCm but provides more control than PyTorch and jax, and hence typically yields more efficient code.
For specific applications, such as linear algebra or discrete Fourier transform, also highly optimized libraries exist. These provide very high performance while, through the targeted use case, also allow for much simpler usage than, e.g., CUDA. In our work, we will build on a linear algebra formulation of our adaptive PDE solvers. This allows us to use the cuBLAS [cuBLAS] and cuSPARSE libraries [cuSPARSE] that provide high performance while being simple to use with a software interface that is roughly comparable to those of the conventional BLAS library. Furthermore, cuBLAS and cuSPARSE can still be combined with native CUDA code when necessary, i.e. when an operation cannot be expressed efficiently in linear algebra.
@@ 4 CUDA multigrid
In this section we discuss our extension of the adaptive finite element library Gascoigne 3d to use GPU accelerators for a wide range of computations. The extensions have been designed with two use cases in mind: first, non-stationary but linear systems of partial differential equations, where a single system matrix is reused in every time step or at least in many time steps, and, second, nonlinear problems that can be formulated in terms of fixed matrices that do not need re-assembly. The second case includes a pressure-projection based solvers for the nonstationary nonlinear Navier-Stokes equations but can be extended to more complex flow problems such as the Boussinesq approximation [PAMMBoussinesq]. Both problem classes have in common that matrices are assembled only once and the complete workflow can be formulated in terms of matrix vector products and the solution of linear systems of equations. Algorithms 2 and 3 show the typical workflow for the problem types. Computations highlighted in blue are completely performed on the GPU, orange marks transfer of data between GPU and CPU.
@@ 4 Assemble Right Hand Side;
It is possible to also port the computations of general nonlinear problems to a GPU. However, keeping the flexibility and support of different discretizations and triangulations is not as easily handled for the ill-structured setup of typical finite element assemblies of system matrices and residuals in the general nonlinear case. We hence defer this to future work.
@@ 4.1 General architecture
Our principal approach for the CUDA parallelization of Gascoigne 3d is to formulate all operations in terms of (potentially sparse) matrix-vector and matrix-matrix products and to realize these with the highly optimized cuBLAS [cuBLAS] and cuSPARSE [cuSPARSE] libraries. This approach differs from the existing implementation in Gascoigne 3d where many operations such as the handling of hanging nodes or the multigrid mesh transfer are based on the local connectivity of the degrees of freedom as function of mesh and discretization. However, this functionality can easily be realized as matrix-vector products, including for the handling of hanging nodes and all mesh transfer operations. This is, e.g., the approach taken from the outset in the deal.II finite element library [dealII].
All high level operations in Gascoigne 3d are performed on abstract interface classes to vectors and matrices to allow for MPI parallelization [P1, P2]. This has been retained in the CUDA-version so that the high-level control flow and complex algorithms (such as Newton, multigrid, GMRES) are unchanged since they only operate on the interfaces. The approach minimizes the changes that were overall required and allows to flexibly retain both the CPU and the GPU backend.
@@ 4.1.1 Storage
Gascoigne 3d uses a block-wise memory concept in which the various components of a partial differential equation (e.g. the 3 deformation unknowns in 3d solid mechanics or the pressure and the three velocity components in fluid mechanics) are stored next to each other. Matrices are double-indexed and the outer index $(i,j)$ refers to the mesh node yielding the entry $a_{i,j}\in\mathbb{R}^{n_{c}\times n_{c}}$ that stores (aligned in memory) the local couplings between the solution components as a matrix. In general, this setup helps to efficiently use caching [BraackRichter2006d, P1]. Using cuSPARSE this block-matrix format is not supported. While vectors are directly transferred to the GPU using the same memory layout, the inner matrix blocks must be resolved and matrices are stored in the usual CSR format. Fig. 2 shows a visualization of the matrix storage concept on the CPU (left) and the GPU (right).
@@ 4.1.4 Geometric Multigrid CUDA
As a concrete example, we describe in the following how geometric multigrid algorithm in Algo. 1 is realized in CUDA-Gascoigne-3d. All computations are entirely performed on the GPU. Hence, data transfer overhead is incurred only at the beginning and end of the computations, see also Algorithms 2 and 3.
@@ 4.1.5 GMRES as linear solver on the GPU
For reasons of numerical stability, e.g. on non-uniform meshes, the multigrid iteration can often not be used directly as linear solver. They are then employed as preconditioner in a GMRES iteration. Our GMRES solver follows the approach described in [Saad1996, Section 6.5.3], using the modified Gram Schmidt algorithm for orthogonalization and Givens rotation for solving the resulting overdetermined linear system. Due to the multigrid preconditioning, we avoid stability issues in the Gram Schmidt orthogonalization and never have to perform more than 5-10 GMRES steps. The Gram Schmidt iteration is also completely run on the GPU using cuSPARSE and cuBLAS operations. As we use a highly efficient geometric multigrid solver as preconditioner, a maximum of $n_{G}\leq 10$ GMRES steps is usually required. Hence, the resulting overdetermined system for finding the GMRES solution is very small and its step-wise transformation to a diagonal matrix with Givens rotations is performed on the CPU. The cost for transferring these $n_{G}^{2}$ entries between GPU and CPU is small. Details on GMRES are given in [Saad1996].
@@ 5 Results
All CPU computations are carried out on an AMD EPYC 7773X. For the relatively small problem sizes we consider, our scaling test in Fig. 3 shows that there is only a minor benefit from using a large number of threads. The calculations are dominated by sparse matrix-vector multiplications and the efficiency is therefore limited by the memory bandwidth. This has already been shown in a detailed parallel analysis of fluid-structure interactions based on Gascoigne [Failer2020]. A transition to matrix-free finite elements—see, e.g., [Kronbichler2019]—would thus offer the greatest potential for improving parallel efficiency. However, this would require a fundamental redesign of the current implementation. We therefore limit the OpenMP based parallel code to 8 threads for the comparisons. The GPU computations are performed on the same machine using an NVIDIA H100 PCIe card with 80 GB of VRAM. When using the GPU to accelerate the multigrid solver, those parts of Gascoigne that reside on the CPU still run on 8 parallel threads. We always indicate wall-clock times and separately specify the different contributions of Algorithms 2 and 3.
@@ 5.1 Transport-diffusion equation
Figure 4: Transport-diffusion equation: Wall-clock times CPU vs. GPU on the finest mesh level with 1 000 0001\,000\,000 degrees of freedom (left: linear scale, right: log scale).
@@
We consider the backward Euler method with a time-step size $\Delta t=0.02$. Uniform quadrilateral spatial meshes with a range of sizes given by $h=2^{-n}$ where $n\in\{7,\dots,10\}$ are used to demonstrate the scaling efficiency of the GPU accelerated version. Hence, on the coarsest mesh with $h^{-2}=2^{2n}$ elements, the problem has $(2^{7}+1)^{2}=16,641$ degrees of freedom whereas the finest mesh comprises $(2^{10}+1)^{2}\approx 1,000,000$ degrees of freedom. In each of the $T/\Delta t=100$ time steps, a linear system must be solved. To ensure the robustness of the method, we use the GMRES solver, preconditioned with multigrid, to approximate the linear problems, cf. Sec. 4.1.5. Fig. 4 shows the wall-clock times comparing the CPU and GPU implementations on the finest mesh level. The labels init, rhs and solve refer to the initialization (mainly assembly of system matrix), to the computation of the right hand side, and to the actual solution, respectively. Matrix assembly and the computation of the right hand side involve numerical quadrature over the mesh elements which is not ported to the GPU. Hence, no speedup is observed, GPU timings can even be higher, as copying the matrix to the GPU involves some overhead, see also Fig. 5. The time for solving the linear problems, however, is reduced from $304.75\,\mathrm{s}$ on the CPU to $9.0\,\mathrm{s}$ on the GPU, a factor of about 34; see also Table 1 for the raw data. The geometric multigrid solver is robust on all meshes such that the number of linear steps varies only slightly between the mesh levels. For large problems, the non-accelerated assembly of the right hand side rhs hence becomes dominant. In Section 5.3 we describe how this term can be efficiently transferred to the GPU by using mass lumping.
@@ 5.2 Linear elasticity
For time discretization of (6)-(7) the backward Euler method is used with a step size of $\Delta t=0.025$ resulting in $2.5/\Delta t=100$ time steps.
@@ 5.3 Navier-Stokes equations
$(\mbox{div}\,\boldsymbol{u},\xi)\;=\;$ $0\qquad$ $\forall\xi\in Q.$ (11)
@@
$\frac{1}{k}(\boldsymbol{u}_{h}^{m},\boldsymbol{\chi})_{*}$ $\;=\;(\boldsymbol{f}^{m-1}+\frac{1}{k}\boldsymbol{u}^{m-1}_{h},\boldsymbol{\chi})_{*}-A((\boldsymbol{u}^{m-1}_{h},p^{m-1}_{h}+q^{m-1}_{h});(\boldsymbol{\chi},0))\qquad\forall\boldsymbol{\chi}\in\boldsymbol{V}_{h}.$
@@
$(\nabla q^{m}_{h},\nabla\varphi)$ $\;=\;-\frac{1}{k}(\mbox{div}\,\boldsymbol{u}_{h}^{m},\varphi)\qquad\forall\varphi\in S_{h}.$
@@
$(p_{h}^{m},\varphi)_{*}$ $\;=\;(p_{h}^{m-1}+q_{h}^{m},\varphi)_{*}-(\nu\,\mbox{div}\,\,\boldsymbol{u}_{h}^{m},\varphi)\qquad\forall\varphi\in Q_{h}.$
@@
Moreover, we apply a mass lumping approach also in the nonlinear term, which enables us to implement the convective term as a matrix-vector product with a pre-assembled sparse matrix, see also [KR2024]. In particular, let $\mathcal{I}_{h}:\boldsymbol{V}\to\boldsymbol{V}_{h}$ be the nodal interpolation operator. For each $\boldsymbol{v}_{h},\;\boldsymbol{\chi}_{h}\in\boldsymbol{V}_{h}$, we approximate
@@
The error caused by mass lumping in the zeroth-order term (12) and the convective term (13) are of the same order as the polynomial approximation error and therefore do not affect the error asymptotics. This approach is sometimes denoted as ‘fully practical finite element method’ [Barrett2001] as it allows for error estimates that reflect the full error including numerical quadrature. We have recently published the complete error analysis for an explicit pressure correction method based on this approximation [KR2024]. Both theoretical analysis and numerical examples show that the explicit treatment based on the interpolation of the nonlinearity does not lead to any additional error.
@@
$\big(\boldsymbol{u}_{h}\otimes\boldsymbol{u}_{j},\nabla\boldsymbol{\chi}\big)_{i=1}^{N_{nodes}}\approx-C_{h,x}\boldsymbol{v}^{1}-C_{h,x}\boldsymbol{v}^{2}-C_{h,x}\boldsymbol{v}^{3}.$ (15)
@@ 5.3.1 Explicit projection solver
For justification and well-posedness of this equation, see [rannacher1982]. Moreover, note that the velocity solutions $\boldsymbol{u}_{h}^{m}$ are so-called predictor solutions. The corrector velocities were eliminated as described in [guermond1996]. The complete error analysis for this fully explicit variant of a pressure correction scheme is found in [KR2024] and its extension to the Boussinesq approximation is described in [RKR2024].
@@ 5.3.3 Numerical Results
Memory is only transferred for initialization (first copy of matrices and vectors) as well as for the GMRES solver. Here however, only single floats, e.g. results of scalar products or short matrices used in the GMRES orthogonalization must be copied. Their dimension is in the order of number of GMRES steps (always less than 10). By combining all optimizations the overall runtime for the $8/\Delta t=40\,000$ time steps is reduced from about $26\,800\,\mathrm{s}$ (about $7.5\,\mathrm{h}$) using 8 CPU cores to about 1 700 s (about $30\,\mathrm{min}$) with the GPU, i.e. we obtain a speedup by a factor of approximately 16.
@@ 5.4 Limitations
For speed, the bottleneck of the GPU implementation is the memory bandwidth. The increase in bandwidth by a factor of 10 from the DDR4 RAM accessed by the CPU to the HBMe2 memory utilized on the H100 GPU (2 TB/s) is complemented by a similar increase in potential compute throughput. An analysis using NVIDIA’s Nsight profiler shows that roughly $2/3$ of the GPU cycles are spent on sparse matrix-vector products, which have an arithmetic intensity of 0.39FLOP/byte (0.15FLOP/byte for the transposed variant) and therefore achieve just $\frac{1}{32}$ ($\frac{1}{83}$) of the peak double-precision performance. As part of cuSPARSE, the kernels are already highly optimized with a memory throughput of $91\%$ ($82\%$) and the low arithmetic intensity is inherent to the computational problem. Major improvements would therefore require fusing some of the roughly 3500 kernel calls needed for a single time-step together in order to reuse the data. This is not practical with the current approach that heavily leverages the cuSPARSE library. All the custom kernels together use less than $3\%$ of the GPU time. While some of these could be eliminated by assembling rectangular block matrices in Gascoigne 3d, the speedup would therefore be negligible. The most promising direction for further optimizations is therefore a switch to single precision, since this effectively doubles the arithmetic intensity.